# Juru: Legal Brazilian Large Language Model from Reputable Sources

### v1

@@ 5.3 Data Contamination
While positive results are evident, there’s a noticeable risk of contamination in the pretraining dataset.This risk arises from web scraping pretraining data in the latter half of $2023$, after the publication of all exams used. Although the origin of data from reputable sources in Brazilian academic works likely doesn’t include the data from the law benchmark, this risk of validity persists. Hence, to mitigate data contamination, the continuation of this work involves annotating new exams released in 2024 and 2025.
6 Conclusion
@@
In this study, we presented the Juru model, a large language model specialized for the Brazilian legal domain using $3.7$ billion tokens. Despite its smaller size and limited data, we observed an expressive improvement in the model’s ability to resolve multiple-choice questions within the Brazilian legal domain. Our primary finding highlights the effectiveness of domain specialization in improving LLM performance with fewer data, thus reducing pretraining costs, at the expense of the model’s performance degradation in other knowledge domains within the target language.
For future work, our primary aim is to incorporate evaluations of the Juru model with new benchmarks that fall outside the model’s knowledge cut-off to mitigate the possibility of data contamination. Additionally, it’s pertinent to explore methods for mitigating the observed degradation in domains outside the model’s specialization.
7 Acknowledgements

### v2

@@ 6 Limitations
While positive results are evident, there’s a possible risk of contamination in the pretraining dataset. This risk arises from web scraping pretraining data in the latter half of 202320232023, after the publication of the OAB-2023 and MMLU exams. Although the dataset was primarily sourced from reputable Brazilian academic and legal documents, and it is unlikely to contain the test questions, the risk of indirect contamination cannot be fully ruled out. To address this, we plan to include newly released exams from 2025 and 2026 in future evaluations.
All experiments reported in this paper are based on a specific subset of Brazilian legal data. Whether the observed trends generalize to other subdomains or other types of data than those originated from reputable sources remains an open question. Exploring a broader range of legal corpora, including news articles, blogs, and jurisprudential commentaries, may reveal different specialization behaviors.
Our work is subject to the same validity challenges commonly faced in the evaluation of LLMs, as noted in prior studies [8, 10, 11]. In the Brazilian context, what constitutes legal proficiency? Is it the ability to recall legislation, generate legally sound summaries, or predict judicial decisions? We argue that these skills are interrelated. Therefore, improving performance on one legal task may correlate with improvements in others, as they often rely on overlapping capabilities. We do not claim that Juru excels across all legal language processing tasks. Rather, we use performance on Brazilian legal standardized exams, designed to assess human legal competence, as a practical proxy for specialization effectiveness. The observed improvements on these benchmarks suggest that domain specialization may transfer to a broader range of legal tasks.
7 Conclusions and Future Work
@@
In this study, we introduced Juru, the first LLM specialized in the Brazilian legal domain, pretrained on 1.91.91.9 billion unique tokens. Despite its smaller size and limited data, we observed an expressive improvement in the model’s ability to resolve multiple-choice questions within the Brazilian legal domain. Our findings confirm that domain specialization can be an efficient strategy for improving LLM performance at a lower compute cost. However, this comes with a trade-off: diminished performance on tasks unrelated to the target domain.
Additionally, the comparison between Portuguese and English general test suites revealed a more pronounced decline in the English benchmarks. This suggests that the extent of forgetting may be linked to domain similarity: the lower the similarity the previously acquired knowledge has to the specialized domain, the higher the risk of forgetting.
For future work, our primary aim is to incorporate evaluations of the Juru model with new benchmarks that fall outside the model’s knowledge cut-off to mitigate the possibility of data contamination. A promising direction is to investigate whether aggregating domains with high similarity during continued pretraining could enhance specialization performance without catastrophic forgetting of previously learned knowledge. Deeper understanding of these dynamics may lead to the development of LLMs with less compute through targeted domain specialization. We also release all Juru checkpoints from the experiments presented in this paper to further support future research on these topics, particularly within the Brazilian legal domain.
7.0.1 Acknowledgements
@@
[…]
The authors declare no conflicts of interest related to the content of this article.